OpenStudy (anonymous):

Multiply and simplify: (4z - 3)(z - 1)

OpenStudy (anonymous):

foil

OpenStudy (anonymous):

use the distributive property

OpenStudy (abb0t):

4z(z) - 4z - 3z +3

OpenStudy (anonymous):

\[ (4z - 3)(z - 1)=4z+\times z-1\times 4z-3\times z+3\times 1\]

OpenStudy (anonymous):

combine like terms, get \[4z^2-7z+3\]

OpenStudy (tkhunny):

\((x+1)(x-2)\) -- \(FOIL\) \((x^{2} + x + 1)(x-2)\) -- \(FOI_{1}I_{2}E_{ish}L\) This is why FOIL is silly. It has limited application and cannot be generalized. Simply learn to multiply. It is EXACTLY like the original hand algorithm for long numbers.